OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost?

OpenStudy (anonymous):

x = cost of 1 day in hotel y = cost of 1 day at the amusement park the equation for plan A is: 3x + 2y = 450 the equation for plan B is: 5x + 3y = 700 then solving for x: -3 * (3x + 2y = 450) = -9x - 6y = -1350 2 * (5x + 3y = 700) = 10x + 6y = 1400 the y's cancel out leaving you with x = 50 so option 1 is the answer

OpenStudy (anonymous):

ok thanks